OpenStudy (anonymous):

Find the radius, diameter, and area of a circle with circumference 6π. How do I even begin to do this problem?

jhonyy9 (jhonyy9):

r=3

OpenStudy (anonymous):

circumference =2*pi*r=6*pi

jhonyy9 (jhonyy9):

D=6

OpenStudy (anonymous):

Thank you!